OpenStudy (anonymous):

how to do working for square root -3?

OpenStudy (anonymous):

this is not a real number If you want further it should take as\[\sqrt{-1}\times \sqrt{3}\]

OpenStudy (anonymous):

1.73205081 i , not sure how to work it out.

OpenStudy (anonymous):

the\[\sqrt{-1}=i\] it represents imaginary numbers not real

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

correct for??

OpenStudy (anonymous):

1.73205081 i It's an imaginary answer because it represents a number that is not real.